Each Tuesday, Ryan Airlines reduces its 1-way ticket from Fort Wayne to Chicago from $120 to $30. To receive this special $30 price, the customer must buy a round-trip ticket. Ryan has a nonrefundable 25% penalty fare for CX; it estimates that about eight-tenths of 1% will CXtheir reservations. The airline also estimates this special price will cause a passenger traffic increase from 400-800. Ryan expects revenue for the year to be 53.3% higher than the previous year. Last year, Ryan’s sales were $489,000. To receive the special rate, Janice Miller bought 2round-trip tickets. On other airlines, Janice has paid $95 round trip (w/no cancellation penalty).

a. Calc the % discount Ryan is offering. (nearest 100th %.)

b. Calc the % passenger travel will increase.

c. Calc the sales for new year.

d. Calc Janice’s loss if she cancels one round-trip flight.

e. Approx how many more cancellations can Ryan Airlines expect (after Janice’s cx)? (whole #.) 

1
Expert's answer
2021-08-02T17:22:54-0400

a) price of one round trip before discount ="2\times 120=\$240"

price of one round trip after discount=120+30=150

percentage discount offered "=\frac{240-150}{240}\times 100\%=90/240\times 100\%=37.5\%"


b) percentage increase in discount"=\frac{800-400}{400}\times100\%=100\%"


c) 1% of 800=80

Since the number of passengers estimated to cancel their reservations "=8/10\times 8=6.4" which is approximately 6

Then, sale for new year=25% penalty of 6 people+794 reservations

"=25/100\times 6\times 150+794\times 150=225+119100=\$119325"

d) cost of one round trip ="\$150"

so, Janice loss =25% cancellation penalty"=25/100\times 150=\$37.5"

e) Since the total estimate of cancellation is 6, Ryan airlines can expect 5 more cancellation

Number of cancellation =5 more people